plese help
a long distance trucker traveled 224 miles in one direction during a snow strom. the return trip in rain weather was accomplished at double the speed and took 4 hours less time. find the speed going
thank you for your time

Answer by stanbon(75887)

a long distance trucker traveled 224 miles in one direction during a snow strom.
the return trip in rain weather was accomplished at double the speed and took 4 hours less time.
find the speed going
---
Going DATA:
distance = 224 miles ; rate = x mph ; time = 224/x hrs.
---------------------------
Return DATA:
distance = 224 miles ; rate = 2x mph ; time = 224/(2x) hrs
----
Equation:
Go time - Return time = 4 hrs
224/x - 224/(2x) = 4 hrs
----
Multiply thru by 2x to get:
2*224 - 224 = 8x
224 = 8x
x = 28 mph (Going rate)
=============================
Cheers,
Stan H.